ganeshie8 (ganeshie8):

\[\sigma(n+1)-\sigma(n-1)=2\]

Parth (parthkohli):

twin primes?

OpenStudy (sparrow2):

what is a question? :)

Parth (parthkohli):

idk if this is true for anything other than twin primes

ganeshie8 (ganeshie8):

Wow, Is that too easy ha !

OpenStudy (sparrow2):

can you specify more clearly sigma and n?

ganeshie8 (ganeshie8):

\(\sigma(n) = n+1 \iff \text{ n is prime}\)

Parth (parthkohli):

yes, I see that but I'm trying hard to check why it cannot be something of this form\[\sigma(n+1) = n+1 + k\]\[\sigma(n-1) = n-1 + k\]that is to say that the proper divisors have the same sum. let me think.

ganeshie8 (ganeshie8):

Interesting...

OpenStudy (imdwerty):

:/ what we need to do in this question tho?..

ganeshie8 (ganeshie8):

@sparrow2 \(\sigma\) is the sum of divisors function http://mathschallenge.net/library/number/sum_of_divisors

OpenStudy (imdwerty):

sigma function is defined for positive integers and its minimum value is 1 and for \(\sigma(x)=1\) ; x=1 so for \(\sigma(n-1)+\sigma(n+1)=2\) both (n-1) and (n+1) must be 1 and this is not possible

Parth (parthkohli):

you answered a different question lol

OpenStudy (imdwerty):

which question are we solving?

OpenStudy (sparrow2):

so we are trying to prove that sigma(n)=n+1 is working only for primes?

Parth (parthkohli):

ok I will write a quick program to find non-primes with the same property

ganeshie8 (ganeshie8):

@ParthKohli 435 is the first counter example it seems

Parth (parthkohli):

oh you already did it. great :)

OpenStudy (kainui):

Are there infinitely many counter examples? Are there nontwin primes that satisfy the equation AND also satisfy this equation? \[\tau(n+1)=\tau(n-1)\]

OpenStudy (sparrow2):

if you plug n=4 : sigma(5)-sigma(3)=6-4=2.. is 4 also solution?

OpenStudy (kainui):

Yeah, it definitely is @sparrow2 :D

OpenStudy (sparrow2):

then every n,there n+1 and n-1 are primes are solutions too,there are many solution

OpenStudy (kainui):

Yeah, when \(n+1\) and \(n-1\) are primes, these are called twin primes, so you just found 3 and 5 are twin primes, another pair of twin primes is 11 and 13.

OpenStudy (sparrow2):

less then 1000 there are 36 solutions :D

OpenStudy (sparrow2):

if my program works well :D i'm sure it works

OpenStudy (sparrow2):

are twin primes infinite?

OpenStudy (sparrow2):

oh ist's unsolved :D

OpenStudy (samigupta8):

Ok so well you might have found it yourself because that formula works well for prime numbers only....